OpenStudy (anonymous):

Find the inverse. y=2 square root 5 x^3+5

OpenStudy (anonymous):

\[y=2\sqrt[5]{x^3+5}\]

OpenStudy (anonymous):

I thought you were supposed to switch x and y first?

OpenStudy (anonymous):

Fair enough. We are used to doing it differently but the results are the same. So switch x & y. Then you'll need to isolate y.

OpenStudy (jhannybean):

First checkif this function passes the horizontal line test for all values of x, and whether or not it is 1 to 1.

OpenStudy (anonymous):

\[x/2=\sqrt[5]{y^3+5}\]

OpenStudy (anonymous):

Great. Now to get rid of the 5th root. What's the inverse operation of taking the 5th root?

OpenStudy (anonymous):

\[(x/2)^5=y^3+5\]

OpenStudy (anonymous):

Terrific. Now get rid of the 5 by subtracting from both sides.

OpenStudy (anonymous):

\[(x/2)^5-5=y^3\]

OpenStudy (anonymous):

Right on. Now solve for y. What's the inverse operation of raising a number to exponent 3?

OpenStudy (anonymous):

\[\sqrt[3](x/2)^5-5=y\]

OpenStudy (anonymous):

Excellent. Don't forget to take the cube root of everything on the left hand side. Well done.

OpenStudy (anonymous):

\[\sqrt[3]{\left( \frac{ x }{ 2 } \right)^5-5}=y\]

OpenStudy (anonymous):

Thank you! Do you think you could help me with one more? It's easier than this one.

OpenStudy (anonymous):

I'm terribly sorry. I have to run right now. But there's lots of great people on here to help you. Good. luck.